### v1

# Prompt to Pwn: Automated Exploit Generation for Smart Contracts

Abstract
We explore the feasibility of using LLMs for Automated Exploit Generation (AEG) against vulnerable smart contracts. We present ReX, a framework integrating LLM-based exploit synthesis with the Foundry testing suite, enabling the automated generation and validation of proof-of-concept (PoC) exploits. We evaluate five state-of-the-art LLMs (GPT-4.1, Gemini 2.5 Pro, Claude Opus 4, DeepSeek, and Qwen3 Plus) on both synthetic benchmarks and real-world smart contracts affected by known high-impact exploits. Our results show that modern LLMs can reliably generate functional PoC exploits for diverse vulnerability types, with success rates reaching up to 92%. Notably, Gemini 2.5 Pro and GPT-4.1 consistently outperform others in both synthetic and real-world scenarios. We further analyze factors influencing AEG effectiveness, including model capabilities, contract structure, and vulnerability types. We also collect the first curated dataset of real-world PoC exploits to support future research.
Introduction
Once smart contracts were deployed, even a small vulnerability may be permanently exploitable, resulting in substantial financial losses. We show an example.
In February 2025, a smart contract exploit on Bybit’s Safe multi-signature wallet allowed attackers to upgrade the contract implementation and drain 1.5 billion US dollars (Rommen and Sehmbi 2025).
To address the security issues, many vulnerability detection tools have been proposed, including Slither (Feist, Grieco, and Groce 2019), Mythril (Nikolić et al. 2018), and Oyente (Luu et al. 2016), but existing static and symbolic analysis tools suffer from low accuracy and limited scalability and perform poorly in practice (Sendner et al. 2024).
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of applications, including code-related tasks such as generation (Laurie et al. 2022), summarization, and bug fixing (Pearce et al. 2022). Not surprisingly, LLMs also hold strong potential for assisting in the identification of smart contract vulnerabilities (Chen et al. 2025; David et al. 2023; Xiao et al. 2025).
We step further to explore their effectiveness in generating verifiable exploits. We focus on a few research questions:
  • RQ1: How well can LLMs perform automated exploit generation (AEG) for smart contract vulnerabilities on both benchmarks and real-world attack data?
  • RQ2: What factors affect AEG effectiveness: contract properties (e.g., size, complexity), vulnerability types, or specific prompt design?
  • RQ3: What defensive practices can mitigate such LLM-driven threats?
To investigate whether LLMs can move beyond detection to AEG and validate executable proof-of-concept (PoC) exploits for smart contracts, we propose ReX, a new framework that integrates LLM-based exploit synthesis with the Foundry testing environment, enabling end-to-end exploit generation, compilation, execution, and verification. We evaluate five most latest LLMs (GPT-4.1, Gemini 2.5 Pro, Claude Opus 4, DeepSeek-R1, Qwen3 Plus) on their ability to generate full exploit contracts and test scripts, targeting eight distinct real-world vulnerability types (e.g., reentrancy, integer overflows, improper access control).
To further improve output automation, we design three optimization techniques, including prompt refactoring, compiler feedback loops, and templated test harness generation. These enhancements significantly improve success rates for underperforming models. Our results show that:
  • GPT-4.1, Gemini 2.5 Pro, and Claude Opus 4 achieve success rates 80%+ on most vulnerability categories, with GPT-4.1 peaking at 92.5% on arithmetic bugs.
  • DeepSeek-R1 and Qwen3 Plus produce valid exploits in 60–70% of cases but show inconsistency across more complex vulnerability types.
  • Over 76% of generated artifacts pass compilation and test execution end-to-end without human intervention.
In addition, we curate the first dataset111provided upon request. of manually written PoC exploits targeting real-world contracts, consisting of 38+ audited attack cases from expert security researchers.
Finally, we validate LLM-generated exploits on smart contracts that suffered from historical high-impact attacks. In 4 out of 10 cases, Gemini 2.5 Pro was able to independently rediscover valid exploits. It not only demonstrated autonomous exploit discovery capabilities, but also exhibited expert-like reasoning strategies similar to those employed by seasoned smart contract auditors.
To facilitate reproducible research and open science, we provide the code, including the detailed prompts and raw experiment data as supplementary materials, and will open-source them upon paper acceptance.
Technical Warmups
Smart contract vulnerabilities.
Smart contracts are vulnerable to bugs and exploits (Tolmach et al. 2021). Once deployed, contracts cannot be altered, leaving any flaws permanently exposed. To standardize and detect such issues, the SWC Registry (swc 2018) categorizes common vulnerabilities like reentrancy (SWC-107), integer overflows (SWC-101), and access control flaws (SWC-105), forming the basis for many tools and benchmarks. As decentralized applications grow in complexity, attackers increasingly leverage sophisticated techniques such as arbitrage (Torres, Steichen et al. 2019; Cernera et al. 2023), transaction order manipulation (Eskandari, Moosavi, and Clark 2019), and proxy-based upgrade attacks (Meisami and Bodell III 2023).
LLMs.
We evaluate five latest LLMs in this work. GPT-4.1 (OpenAI) is known for strong reasoning and consistent performance across code and language tasks. Gemini 2.5 Pro (Google) handles text, code, and images well, matching GPT-4-class performance. Claude Opus 4 (Anthropic) offers long-context reasoning and is suited for safety-critical tasks. DeepSeek-V2 delivers high performance in bilingual and code tasks. Qwen3 Plus (Alibaba) provides strong code and dialogue capabilities, especially in English and Chinese.
Contract testing suite.
We use Foundry, a Rust-based toolchain for smart contract development and testing (Fou 2025). Foundry supports dependency management, compilation, deployment, and testing. Its built-in test runner (forge test) enables efficient, large-scale validation of Solidity exploits, outperforming tools like Hardhat and Truffle.
Static analytic tools.
We use two. SUMO is a static analysis tool for smart contracts that estimates code understandability through a composite complexity score (Barboni, Morichetta, and Polini 2022). It captures Solidity-specific features like control-flow depth, nesting, and branching. Slither is another widely used static analysis tool for detecting vulnerabilities and computing code metrics in Solidity contracts (Feist, Grieco, and Groce 2019). It provides accurate and fine-grained structural insights.
ReX and Datasets
Figure 1: Overview of the ReX framework
As shown in Figure 1, ReX is a fully automated pipeline and proceeds in five steps:
Step 1: data preprocessing. We first clean the input smart contracts by removing comments and non-functional content to eliminate noise and avoid misleading the LLMs. We ensure that the LLM focuses only on the core contract logic when generating the exploit.
Step 2: script generation. Gven a vulnerable smart contract, through carefully designed prompt, the LLM will generate two related Foundry scripts: the first is an exploit contract designed to exercise the vulnerable code path; while the second is a test contract that validates whether the exploit succeeds under some specific conditions.
We enabled the LLMs to iteratively optimize its prompts (Chen et al. 2025; Grubisic et al. 2024; Sepidband et al. 2025). As a result, it is capable of generating exploit and test scripts with intact import paths, designed to be as compilable and executable as possible. At the same time, we enabled the model’s reasoning mode by guiding it to reason step by step, aiming to improve the accuracy of its responses.
Step 3: optional script optimization. Initial experiments reveal that many LLM outputs suffer from minor but recurring issues that prevent compilation. To address this, we include an optional postprocessing step that automatically fixes common errors:
  • EIP-55 address checksum: Non-checksummed Ethereum addresses are normalized to conform with EIP-55.
  • Missing payable casts: Calls to value-transferring functions often lack the required payable cast, which we insert automatically.
Step 4: compilation and testing. The generated scripts are integrated into a Foundry project and passed through the following toolchain:
  • forge init sets up the testing environment;
  • forge build compiles the contracts;
  • forge test -vvvv runs the tests with detailed output, including gas usage, logs, and traces.
This step verifies the scripts both syntactically and semantically within a local Ethereum-like environment.
Step 5: Iterative feedback loop If compilation or testing fails, the error messages and scripts are returned to the LLM, which attempts to correct and regenerate them. This loop repeats until a valid exploit is found or a maximum retry limit is reached. This feedback-driven refinement is inspired by prior success in LLM-based program synthesis (Bi et al. 2024; Grubisic et al. 2024), and significantly boosts the success rate of exploit generation.
What datasets used for evaluation? We employ two datasets to evaluate ReX’s ability: one well-established benchmark for controlled testing (Durieux et al. 2020), and one curated collection of real-world attack cases.
SmartBugs-Curated. The dataset was selected for the following reasons:
  • As a classical smart contract vulnerability dataset. It has been widely used in both smart contract vulnerability detection community and related researches (Ferreira et al. 2020)(Chen et al. 2025)(Durieux et al. 2020);
  • All contracts are carefully labeled with their respective vulnerabilities and have been validated over time, contributing to the robustness and credibility of the dataset;
  • The simplicity and clarity of the dataset make it particularly suitable for the initial evaluation of LLM’s capability to generate vulnerability exploitation PoCs. Smartbugs-curated provides a collection of vulnerable Solidity smart contracts organized according to the DASP taxonomy (Chen et al. 2025).
Web3-AEG. To assess LLM performance on realistic attack scenarios, we construct a new dataset named Web3-AEG, comprising historical high-impact vulnerabilities from real-world smart contracts. This dataset includes:
  • A collection of publicly disclosed vulnerable contracts exploited from 2021 to 2025, with verifiable source code.
  • Ground-truth PoC exploits manually written and published by professional auditors and white-hat hackers.
To the best of our knowledge, Web3-AEG is the first dataset that enables reproducible, automated evaluation of LLM-generated exploits in real-world settings. It provides a critical benchmark for analyzing the practical utility and generalization ability of LLM-based exploit generators.
Evaluating AEGs
Exp1: Benchmarking on SmartBugs-Curated.
The first experiment (Exp1) provides a baseline assessment to evaluate LLM’s performance in AEG with the structurally simple SmartBugs-Curated dataset.
Since Foundry only stably supports compiling smart contracts with Solidity version 0.8 and above, we first perform version migration by using LLMs to rewrite all contracts to use version 0.8.26. For arithmetic vulnerabilities, we explicitly use unchecked blocks to disable overflow and underflow checks. Although Solidity has enabled overflow/underflow checks by default since version 0.8.0, many contracts currently deployed on-chain still use pre-0.8.0 versions.
We manually check the consistency of each smart contract before and after version migration, and exclude contracts that do not meet the criteria. In fact, we found that the vast majority of smart contracts only required updating the compiler version. With version migration, these contracts also differ from their original forms, which helps reduce the likelihood that they appear in the LLM’s training data.
Next, we use the previously mentioned evaluation framework to generate and assess AEG.
For specific vulnerability types, we do not only rely on Foundry test results for evaluation. LLMs may generate exploits that trigger overflows or unexpected reverts-especially in denial-of-service and reentrancy cases—causing the tests to fail. Despite of it, these cases still reflect successful and valid exploit attempts.
Table 1: AEG Performance for Smart Contract Vulnerabilities against of five mainstream LLMs
Vulnerability Gemini 2.5 Pro GPT-4.1 Claude Opus 4 DeepSeek Qwen3-plus
Reentrancy 18/30 (60.0%) 18/30 (60.0%) 19/30 (63.3%) 10/30 (33.3%) 6/30(20.0%)
[1pt/1pt] Access Control 10/18 (55.6%) 9/18 (50.0%) 10/18 (55.6%) 9/18 (50.0%) 4/18(22.2%)
[1pt/1pt] Arithmetic 13/14 (92.9%) 12/14 (85.7%) 12/14 (85.7%) 11/14 (78.6%) 5/14(35.7%)
[1pt/1pt] Bad Randomness 5/7 (71.4%) 4/7 (57.1%) 4/7 (57.1%) 1/7 (14.3%) 1/7 (14.3%)
[1pt/1pt] Front Running 3/4 (75.0%) 1/4 (25.0%) 1/4 (25.0%) 2/4( 50.0%) 1/4(25.0%)
[1pt/1pt] DoS 4/6 (66.7%) 4/6 (66.7%) 6/6 (100.0%) 3/6 (50.0%) 2/6 (33.3%)
[1pt/1pt] Time Manipulation 3/5 (60.0%) 4/5( 80.0%) 4/5 (80.0%) 3/5 (60.0%) 2/5(40.0%)
[1pt/1pt] Unchecked Low Level Calls 17/30 (56.7%) 12/30 (40.0%) 12/30 (40.0%) 15/30 (50.0%) 12/30 (40.0%)
Average Success Rate 67.3% 58.1% 63.3% 48.3% 28.8%
The results are in Table 1. We demonstrated that all LLMs have the capability to perform AEG on smart contract vulnerabilities. Among them, GPT-4.1, Gemini 2.5 Pro, and Claude Opus 4 showed better performance than the others, with Gemini 2.5 Pro achieving the best overall results.
Specifically, Gemini 2.5 Pro achieved the highest average success rate (67.3%) across diverse vulnerability types, excelling in arithmetic (92.9%), front running (75.0%), and unchecked low-level calls (56.7%). GPT-4.1 followed with 58.1%, showing consistent results in arithmetic (85.7%) and time manipulation (80.0%). Claude Opus 4 ranked third (63.3%) but showed strong robustness across all categories, including DoS (100.0%) and access control (85.6%). In contrast, DeepSeek and Qwen3-plus exhibited noticeably lower success rates (48.3% and 28.8%, respectively), struggling especially with complex vulnerabilities such as reentrancy and unchecked low-level calls.
Exp2: Real-world exploits with Web3-AEG.
This experiment (Exp2) evaluates LLMs’ ability to generate valid proof-of-concept (PoC) exploits against contracts that were exploited in high-impact real-world attacks. We use the Web3-AEG dataset, which contains publicly known vulnerable contracts and their corresponding expert-written PoCs. All contracts are tested in their original, unmodified form to preserve authenticity.
We examine two core aspects: (i) determining whether LLMs can generate compilable, functional exploits that demonstrate vulnerabilities; and (ii) analyzing how closely the LLM-generated attack strategy aligns with expert-crafted PoCs. Among the tested models, Gemini 2.5 Pro succeeded in generating four valid PoCs, while GPT-4.1 and Claude Opus 4 each succeeded once.
We show three representative AEG cases.
Case1. Predictable Randomness (RedKeysGame) 0x71e3056aa4985de9f5441f079e6c74454a3c95f0
The contract uses block data as its randomness source, making outcomes predictable. Attackers win every bet by predicting correct numbers off-chain (SlowMist1 2024).
  • LLM PoC. The model analyzes randomNumber(), replicates the logic off-chain, and repeatedly calls playGame() with correct values to ensure wins. The test uses a local simulation to demonstrate exploitability.
  • Expert PoC. The expert forks the BSC chain to a specific block and predicts outcomes using the same reverse-engineered logic. A looped sequence of correct guesses allows the attacker to extract funds at scale.
Both PoCs follow identical attack logic, with the main difference being the use of a local testnet by the LLM versus a mainnet fork by the expert.
Case2. Broken access control (TSURUWrapper) 0x75Ac62EA5D058A7F88f0C3a5F8f73195277c93dA
The contract fails to verify the caller in its handler onERC1155Received, allowing arbitrary minting of ERC20 tokens. This vulnerability resulted in a cumulative losses exceeding 138.78 ETH (SlowMist2 2024).
  • LLM PoC. The attacker contract directly calls the vulnerable function, bypasses the flawed if check, and invokes safeMint() to mint unbacked tokens. Repeating this process enables unlimited token creation.
  • Expert PoC. Human auditor further swaps the minted tokens for WETH, effectively realizing the profit.
The LLM identifies and validates the vulnerability but lacks the DeFi-aware reasoning to chain the exploit into an economic attack, unlike the expert.
Case3. Cross-contract flashloan exploit (Pine) 0x2405913d54fc46eeaf3fb092bfb099f46803872f
The Pine Protocol uses a shared vault for both legacy and upgraded lending pool contracts. This leads to flashloan-based reentrancy exploit (Mutual 2024).
  • LLM PoC. The attacker borrows ETH via a flash loan, triggers a reentrancy callback to make a fake repayment before state update, and drains the vault.
  • Expert PoC. The attack spans multiple contracts. An attacker uses WETH from flashloan to repay debt in the new pool, retrieves NFT collateral, and then repays the old pool by same funds, exploiting the shared vault.
Both exploits use flashloans and reentrancy. LLM follows a single-contract path, directly draining funds from a single contract. The expert leverages cross-contract state inconsistency, showcasing a more advanced attack chain.
Answer to RQ1: can LLMs perform AEGs?
We conclude that LLMs are capable of generating valid exploit PoCs for smart contract vulnerabilities with a relatively high success rate and operational efficiency. Among the five evaluated models, Gemini 2.5 Pro exhibited the strongest performance, successfully producing four working exploits, including those against real-world contracts.
However, current LLMs predominantly generate single-contract exploits. These attacks are typically constrained to vulnerabilities that manifest within the local logic of one contract. In contrast, human experts demonstrate a broader capability to craft complex exploit chains that span multiple contracts, exploit inter-contract state inconsistencies, and interact with DeFi protocols to maximize profit extraction.
Exploring Key Factors Influencing LLM Performance in AEG
We now turn to investigate what determines their success or failure in AEG. Despite strong performance overall, the effectiveness of LLMs is inconsistent across contracts. Some highly complex contracts are easily exploited, while others with simpler structure resist attack. We discuss four factors.
Factor 1: LLM capabilities and failure patterns.
We first hypothesize that the primary determinant of AEG success is the LLM’s inherent capabilities, rather than the target contract structure.
Table 2: Five LLMs on Code Generation Benchmarks
Model Aider LMArena SWE-bench
GPT-4.1 53.0% 1331 55.0%
[1pt/1pt] Gemini 2.5 Pro 83.1% 1496 67.2%
[1pt/1pt] Claude Opus 4 72.0% 1456 79.4%
[1pt/1pt] DeepSeek R1 56.9% 1342 40.6%
[1pt/1pt] Qwen3 61.8% 1291 54.2%
As shown in Table 2, Gemini 2.5 Pro outperforms other models on two widely used programming benchmarks (RankedAGI 2024). These scores correlate with its superior AEG performance in our experiments, reinforcing the view that general coding ability is predictive of LLMs’ exploit generation capacity. Similarly, Claude Opus 4, which also achieved high scores across standard coding benchmarks, demonstrated strong performance in our AEG experiments as well.
Moreover, we identified two recurring failure patterns (P).
  • P1: cryptographic limitations. Many LLMs incorrectly generate non-checksummed Ethereum addresses due to training data biases. LLMs cannot compute Keccak-256 hashes needed for EIP-55 compliance.
  • P2: semantic misunderstandin. LLMs frequently mishandle the ‘payable‘ modifier, omitting necessary casts or using it inconsistently. This reflects a deeper challenge in enforcing compiler-level semantic constraints.
Both success and failure reflect the model’s internal understanding. LLMs function as probabilistic pattern matchers, not reasoning engines. Their systematic errors highlight the boundaries of what current LLMs can achieve in AEG.
Factor 2: properties of target contracts
We next assess whether structural properties of the contracts influence AEG outcomes.
Using the Web3-AEG dataset, we extract source-level metrics such as nSLOC, complexity score, and external call count (Figure 4) with analyses via Cramér's V (Table 4). We also listed main factors of exploited contracts in Table 3.
Table 3: Detailed Analysis
Contract nSLOC Complexity Score
RedKeysCoin_exp.sol 185 113
[1pt/1pt] FIL314_exp.sol 325 225
[1pt/1pt] TSURU_exp.sol 804 527
[1pt/1pt] PineProtocol_exp.sol 817 536
Table 4: Cramér’s V – selected features of Web3-AEG
Feature Web3-AEG Reentrancy
nSLOC 0.233 0.251
[1pt/1pt] Complexity Score 0.248 0.339
[1pt/1pt] ExternalCallsCount 0.095 0.233
[1pt/1pt] InheritanceDepth N/A N/A
[1pt/1pt] HasInlineAssembly N/A N/A
[1pt/1pt] PayableFunc N/A 0.000
The results show weak correlations between structural features and AEG success. This suggests that while complexity may correlate with vulnerability, it is not a reliable predictor of LLM exploitability.
We further validate findings using a reentrancy-specific subset. Similar weak associations reaffirm that surface-level complexity metrics have limited discriminative power. The results have been shown in Figure 3.
Factor 3: Vulnerability Type
AEG success varies by vulnerability class. As shown in Figure 2, arithmetic overflows show the highest success rate in both LLM’s AEG due to their simple structure and fixed pattern. These vulnerabilities lack cross-contract dependencies and are easier for LLMs to detect and exploit.
Figure 2: AEG Success Rate by Vulnerability Type
Factor 4: Prompt Engineering
We experimented with various prompt modifications. Although constraining the output format of the generated contracts showed clear benefits, other prompt modifications resulted in only marginal improvements in AEG. This suggests that LLM performance in AEG is constrained more by internal reasoning capacity than by external instructions.
Answer to RQ2: any factors impact AEG success?
We conclude that the LLM’s internal capacity is the primary determinant of AEG success. Structural metrics such as code length or complexity show only weak correlations. Vulnerability types with predictable structures (e.g., arithmetic overflows) are more exploitable. Prompt optimization has limited effect. Thus, the frontier of AEG performance lies in strengthening LLM reasoning and semantic understanding, not in tweaking prompts or inflating contract complexity.
Defending Against LLM-Based AEG
General suggestions for defense.
Our evaluation of LLM-driven AEG reveals a set of systematic limitations that can be leveraged to design practical defense strategies. Below, we present five defense techniques informed directly by our AEG findings.
(a) with AEG
(b) without AEG
Figure 3: Metrics of selected contracts with and without AEG. (cf. Table 7)
(a) WEB3-AEG 1
(b) WEB3-AEG 2
Figure 4: Contract metrics of WEB3-AEG (cf. Table 8)
  • Externalization via code splitting. We show that successful LLM-generated exploits overwhelmingly target single-contract systems. Cross-contract vulnerabilities remain unexploited. This suggests a practical defense: decompose contract logic into modular components (e.g., separating proxies from logic contracts, using delegatecall to distribute attack surfaces). By forcing the model to reason across multiple contracts, this approach increases the difficulty of generating valid exploit paths.
  • Structural, not superficial, complexity. Unlike traditional code obfuscation, structural complexity (e.g., deep inheritance trees, abstract interfaces, polymorphic dispatch) poses challenges to LLMs. These patterns complicate semantic tracing, function resolution, and vulnerability localization, reducing exploit generation success. Our results indicate that increasing structural abstraction is more effective than adding superficial code noise.
  • Breaking canonical signatures. LLMs are particularly effective at detecting well-known patterns, such as arithmetic overflows. To counter this, defenders can diversify vulnerability contexts by embedding redundant logic, applying unconventional naming conventions, or introducing control-flow indirection near vulnerable statements. These strategies disrupt the model’s pattern-matching capabilities and reduce the reliability of PoC generation.
  • Decoy vulnerabilities. Given that LLMs heavily rely on syntax-level cues to infer vulnerabilities, defenders can intentionally introduce false-positive patterns—decoy code fragments resembling canonical vulnerabilities but with no actual exploitability. These decoys can mislead LLMs during generation, increasing the failure rates.
  • Use of edge syntax and low-level features. Our analysis shows that LLMs struggle to process Solidity’s less common constructs, (e.g., try/catch, inline Yul, inline assembly, and raw opcodes). By implementing critical logic using these features, developers can introduce semantic obfuscation that degrades the model’s ability to synthesize valid exploits. These low-level constructs act as natural barriers to automated reasoning.
Evaluating defense effectiveness.
We randomly selected two contracts for each vulnerability type from SmartBugs-Curated dataset used in our previous experiment, specifically choosing those that have been previously confirmed to be successfully exploited by LLMs Gemini 2.5 Pro or GPT-4.1. The selected contracts cover a wide range of structural complexity, ensuring both simple and complex contract architectures are represented.
  • Individual defensive modifications. We begin by assessing the impact of single-factor defensive changes on the success rate of LLM-based automated exploit generation. To isolate the effect of each factor, we modified each contract using only one of the following strategies: splitting logic into separate modules or contracts, increasing structural complexity (e.g., inheritance depth), altering the invocation pattern of vulnerable functions, introducing new functions that contain vulnerabilities, or using less common language constructs.
Table 5: AEG Success Rate of LLMs Under Single-Factor Defensive Changes
LLM Model Increased Complexity Logic Split Pattern Change Decoy vulnerabilities Rare Constructs
Gemini 2.5 87.5% 81.2% 87.5% 75.0% 87.5%
[1pt/1pt] GPT-4.1 87.5% 87.5% 87.5% 68.8% 81.2%
We observe the impact of single-factor defensive is limited (Table 5). In testing, introducing decoy vulnerabilities proves to be one of the more effective defenses. Our findings suggest that when the contract structure is not overly simplistic, adding such decoy functions can reduce AEG success rates and increase the time required by LLMs to generate a successful exploit. Within our evaluation framework, even when an AEG attempt ultimately succeeds, it generally needs three to four iterations for LLM.
  • Combined defensive modifications Based on our previous experiments for the single-factor defense, we developed a strengthened defense strategy by integrating multiple techniques. First, we increased the semantic complexity of smart contracts through logic splitting, vulnerability pattern transformations, and structural enhancements. Second, we inserted multiple decoy vulnerabilities to reduce the likelihood of LLMs correctly identifying real exploit paths. Finally, we hardened critical components of contracts using low-level constructs like inline assembly to increase semantic complexity of execution logic.
Experimental results were recorded in Table 6. Through applying combined defensive measures, the success rate of LLM-based AEG can be significantly reduced. However, we observe that although the two LLMs did not exploit exactly the same types of vulnerabilities, they were both able to successfully generate attacks against hardened contracts containing BR (Bad Randomness) and TM (Time Manipulation) vulnerabilities. Moreover, in all tests involving these vulnerabilities, the contracts were successfully exploited by both two LLMs.
Answer to RQ3: any defensive practices?
Our experiments demonstrate that while individual defensive strategies have limited impact, introducing decoy vulnerabilities emerges as relatively more effective in hindering LLMs. To improve defense, combined defense strategy is able to greatly reduces the success rate of LLM-based AEG. However, current defense mechanisms still exhibit limitations when applied to specific vulnerability types, including bad randomness and time manipulation, which remain susceptible even under strengthened protection strategies.
Table 6: LLMs-AEG-C
LLMs AEG Success Rate AEG-Type
Gemini 2.5 Pro 43.8% BR, TM
[1pt/1pt] GPT-4.1 37.5% BR, TM
Related Work
Traditional smart contract analysis.
Symbolic execution explores program paths using symbolic inputs and is employed by tools like Mythril (Sharma, Sharma et al. 2022), Oyente (Luu et al. 2016), and Manticore (Mossberg et al. 2019), though it struggles with path explosion and environmental modeling. Static analysis tools such as Slither (Feist, Grieco, and Groce 2019) and Securify (Tsankov et al. 2018) analyze code without execution, enabling early vulnerability detection but often producing false positives and failing with complex behaviors. Fuzzing tools like Echidna (Grieco et al. 2020) and ContractFuzzer (Jiang, Liu, and Chan 2018) execute contracts with random inputs to expose unexpected behaviors, yet may miss deep logic bugs. Meanwhile, machine learning (Ressi et al. 2024) learn from labeled data to detect vulnerabilities in unseen code patterns, but require extensive datasets and lack interpretability.
LLMs in smart contracts.
The application of LLMs in the smart contracts has gained increasing attention with the rise of advanced models like GPT-4 and Gemini. Previous works have demonstrated LLMs’ potential in assisting developers with code generation (Liu et al. 2025), bug detection (Sun et al. 2024), and code summarization (Ma et al. 2025) for Solidity-based smart contracts (He et al. 2024).
LLMs for code tasks.
Development in LLMs have significantly improved their capabilities in code tasks, including code generation and completion. Codex(Wang et al. 2024), CodeGen(Nijkamp et al. 2022), and StarCoder(Li et al. 2023) have demonstrated strong performance on standard benchmarks, achieving competitive or even superhuman results in Python and multi-language tasks.
AEG for smart contracts
AEG was a long-standing goal in software security, mainly used in C/C++ (Avgerinos et al. 2014). In recent years, tools like TeEther (Krupp and Rossow 2018) and Echidna (Grieco et al. 2020) began bridging the gap toward exploit generation, but their applications are still limited in specific vulnerability. The rapid progress of LLMs has opened new possibilities for AI-driven AEG in the smart contract security (Wu et al. 2024).
In parallel to our study, Gervais et al. (Gervais and Zhou 2025) present an LLM-based system for automated exploit generation, which equips LLM with an execution-driven agent to autonomously analyze and attack real-world smart contracts on Ethereum and BNB Smart Chain. Their system achieves a 62.96% success rate on VERITE and extracts up to \$8.59m per exploit. We acknowledge such valuable efforts and contributions by communities.
Conclusion
We present ReX to demonstrate that SOTA LLMs, especially Gemini 2.5 Pro and GPT-4.1, can effectively generate automated exploits for vulnerable smart contracts by synthesizing valid PoC artifacts. We find that exploit success is driven primarily by the model’s reasoning and code generation abilities, not by contract size or complexity. We provide our suggested defense modifications. We also contribute the first curated dataset of real-world PoC exploits to the public.